Find the equation of the line that passes through (2,-3) and is perpendicular to the line 2x-y=10
Answer by lwsshak3(11628)

Find the equation of the line that passes through (2,-3) and is perpendicular to the line 2x-y=10
***
Equation of a line: y=mx+b, m=slope, b=y-intercept
2x-y=10
y=2x-10
slope,m=2
Slope of perpendicular line=-1/2 (negative reciprocal of slope of given line)
equation:
y=-x/2+b
solve for b using coordinates of given point (2,-3)
-3=-2/2+b
b=-2
Equation of perpendicular line:
y=-x/2-2